# Tighter Lower Bounds on Aperiodic Ambiguity Function and Their Asymptotic Achievability

### v2

@@ Abstract
This paper presents tighter lower bounds on the maximum aperiodic ambiguity function (AF) magnitude of unimodular sequences under certain delay-Doppler low ambiguity zones (LAZ). These bounds are derived by exploiting the upper and lower bounds on the Frobenius norm of the weighted auto- and cross-AF matrices, with the introduction of two weight vectors associated with the delay and Doppler shifts, respectively. As a second major contribution, we demonstrate that our derived lower bounds are asymptotically achievable with selected Chu sequence sets by analyzing their maximum auto- and cross- AF magnitudes within certain LAZ.
@@ I Introduction
The maximum magnitude of the AF sidelobes critically influences the target detection and delay-Doppler estimation capabilities of sequences, as the AF peaks of weaker targets may be buried by the sidelobes of strong targets [4]. To achieve reliable detection and estimation, the ideal sequence sets are expected to have zero periodic/aperiodic auto-ambiguity function (AAF) and cross-ambiguity function (CAF) values, except for the AAF peak at the origin of the delay-Doppler plane. However, such sequence sets are impossible due to the limited volume of AFs [4]. In practical applications, a feasible approach is to optimize the local AF sidelobes of sequences over a low ambiguity zone (LAZ) [5] in the delay-Doppler plane, since the maximum delay and Doppler values are generally much smaller than the signal duration and bandwidth, respectively [6, 7, 8, 9, 10, 11].
@@
Correlation functions can be considered as a special case of AFs with zero Doppler. In the literature, two important correlation lower bounds are the Welch bound [12] and the Levenshtein bound [13, 14, 15]. These bounds are useful measures of sequence sets operating under static (or quasi-static) wireless environments. In contrast, the upcoming sixth generation (6G) mobile systems need to deal with 1) highly dynamic environments with high mobility of 1000 km/h (or higher) [16, 17] and 2) integrated sensing and communication (ISAC), as one out of the three new key 6G use scenarios, which is evident from the recently finalized IMT-2030 framework [18, 19]. Driven by these two important applications, it is desirable to understand the AF lower bounds in order to measure the optimality of sequence sets deployed over asynchronous mobile channels. Although the periodic AF lower bounds are investigated in [20] and [5], a comprehensive study of the aperiodic AF lower bounds is missing. There is an aperiodic AF lower bound in [5] but it is derived by extending methodologies from the periodic case. Furthermore, the trade-offs between the lower bounds of the aperiodic AAF and CAF are largely untouched, to the best of our knowledge.
@@
From the construction perspective, while sequences with low aperiodic correlation have been investigated with respect to the Welch bound [12] and the Pursley-Sarwate criterion [21] in [22, 23, 24, 25, 26], little is known on systematically constructed sequences having low aperiodic AF sidelobes. Specifically, the asymptotic behavior of Chu sequence pair [27] was studied in [22] and [23] showing that the maximum auto-correlation and cross-correlation magnitudes of the primary Chu sequence and its conjugate can asymptotically beat the Welch bound [12]. This observation demonstrates that the aperiodic Welch bound is asymptotically achievable, highlighting the effectiveness of Chu sequences. However, similar evidence supporting the achievability of the aperiodic AF lower bounds remains absent.
In this paper, compared to the bound in [5, Theorem 4], we introduce tighter aperiodic AF lower bounds for unimodular sequences with certain LAZ. We also establish trade-offs between the lower bounds of aperiodic AAF and CAF of sequences. Motivated by the Levenshtein bound [13], our core idea is to exploit some properties of the weighted aperiodic AF. However, instead of using a single weight vector in the bounding equation, we introduce a pair of weight vectors to study the different impacts of the delay and Doppler shifts, respectively. Moreover, the structural properties of aperiodic AF with zero delay and the last several delays are also utilized in the derivation.
@@
It is shown that our proposed aperiodic AF lower bounds include several aperiodic correlation bounds as special cases with zero Doppler, such as the Welch bound [12], Sarwate Bound [28], Levenshtein bound [13], Peng-Fan bound [29], and Arlery-Tan-Rabaste bound [15]. Additionally, our analysis reveals the optimal weight vector for the Doppler shift, making the derived aperiodic AF bound a function of the weight vector for the delay shift only. Such an optimal Doppler weight vector evenly weights the Doppler bins within the LAZ, aligning with the cyclic nature of the Doppler domain of the aperiodic AF. In contrast, due to the non-cyclic characteristics of the delay domain of the aperiodic AF, certain delay weight vectors lead to tighter aperiodic AF bounds. Consequently, with the optimal Doppler weight vector, the proposed bounds are shown to be tighter than the aforementioned correlation bounds in non-zero Doppler cases.
@@
Additionally, by analyzing the maximum aperiodic AAF and CAF magnitudes of Chu sequences, we show that certain pairs/sets of Chu sequences are asymptotically order-optimal with respect to our proposed AF bounds over certain LAZ. This analysis substantiates the achievability of these bounds. Finally, we demonstrate the tightness of the proposed bounds through asymptotic analysis and numerical examples.
@@ II-A Definitions
$A_{\bm{x}^{m},\bm{x}^{m^{\prime}}}(\tau,\nu)=\left\{\begin{array}[]{ll}\displaystyle\sum_{t=0}^{N-1-\tau}{x}^{m}_{t}{\left({x}^{m^{\prime}}_{t+\tau}\right)}^{\ast}e^{\frac{j2\pi\nu t}{N}},&0\leq\tau\leq N-1;\\ \displaystyle\sum_{t=0}^{N-1+\tau}{x}^{m}_{t-\tau}{\left({x}^{m^{\prime}}_{t}\right)}^{\ast}e^{\frac{j2\pi\nu t}{N}},&1-N\leq\tau\leq-1,\end{array}\right.$ (3)
@@
We define the LAZ in the delay-Doppler plane as
@@ II-B Properties of Aperiodic AF
$\sum_{\nu=0}^{N-1}{\left|A_{\bm{x}^{m},\bm{x}^{m^{\prime}}}(\tau,\nu)\right|}^{2}=$ $\sum_{\nu=0}^{N-1}\sum_{t=0}^{N-1}x_{t}^{m}{\left(x_{t+\tau}^{m^{\prime}}\right)}^{\ast}e^{\frac{j2\pi\nu t}{N}}\sum_{s=0}^{N-1}{\left(x_{s}^{m}\right)}^{\ast}x_{s+\tau}^{m^{\prime}}e^{\frac{-j2\pi\nu s}{N}}$
$=$ $N\sum_{t=0}^{N-1}{\left|x_{t}^{m}\right|}^{2}{\left|x_{t+\tau}^{m^{\prime}}\right|}^{2}$
$=$ $N\left({N-|\tau|}\right).$ (8)
When $m=m^{\prime}$, $\tau=0$, since $A_{\bm{x}}(0,0)=N$, we have
$\sum_{\nu=1}^{N-1}{\left|A_{\bm{x}}(0,\nu)\right|}^{2}=\sum_{\nu=0}^{N-1}{\left|A_{\bm{x}^{m}}(0,\nu)\right|}^{2}-N^{2}=0.$ (9)
Therefore, for any $\nu>0$, ${\left|A_{\bm{x}}(0,\nu)\right|}^{2}=0$, leading to the conclusion that $A_{\bm{x}}(0,\nu)=0$. Based on $A_{\mathbf{x}}(\tau,\nu)={A_{\mathbf{x}}(-\tau,-\nu)^{\ast}}$, we obtain (7). ∎
Lemma 1 also holds for periodic AF [5]. However, (II-B) reveals some unique properties of aperiodic AF that are further elaborated in Corollary 1.
Another property of the aperiodic AF can be derived based on the Cauchy-Schwarz inequality. Let $d=N-|\tau|$, $d\in[1,N]$, the $d$-th last delay of the aperiodic AF satisfies:
@@
${\left|A_{\bm{x}^{m},\bm{x}^{m^{\prime}}}(\tau,\nu)\right|}^{2}\leq{d^{2}},\quad\forall\,\,|\tau|,|\nu|\in[0,N-1].$ (10)
Proof:
For $10\leq\tau\leq N-$ 1, we have
${\left|A_{\bm{x}^{m},\bm{x}^{m^{\prime}}}(\tau,\nu)\right|}^{2}=$ ${\left|\sum_{t=0}^{N-1-\tau}x_{t}^{m}{\left(x_{t+\tau}^{m^{\prime}}\right)}^{\ast}e^{\frac{j2\pi\nu t}{N}}\right|}^{2}$
$\leq$ $\left(\sum_{t=0}^{N-1-\tau}x_{t}^{m}e^{\frac{j2\pi\nu t}{N}}{\left(x_{t}^{m}e^{\frac{j2\pi\nu t}{N}}\right)}^{\ast}\right)\left(\sum_{t=0}^{N-1-\tau}x_{t+\tau}^{m^{\prime}}{x_{t+\tau}^{m^{\prime}}}^{\ast}\right)$
$=$ $(N-\tau)^{2}.$ (11)
Similarly, ${\left|A_{\bm{x}^{m},\bm{x}^{m^{\prime}}}(\tau,\nu)\right|}^{2}\leq(N+\tau)^{2}$ for $N-1\leq\tau\leq 0$. ∎
@@ III-A Aperiodic AF Lower Bounds Associated to the LAZ
$\mathbf{U}=\begin{bmatrix}{\mathbf{U}^{1}}\\ {\mathbf{U}^{2}}\\ \vdots\\ {\mathbf{U}^{M}}\end{bmatrix},\;\text{with}\;\mathbf{U}^{m}=\begin{bmatrix}{\bm{u}(\bm{x}^{m,0})}\\ {\bm{u}(\bm{x}^{m,1})}\\ \vdots\\ {\bm{u}(\bm{x}^{m,Z_{y}-1})}\end{bmatrix}\;\text{for}\;m=1,2,\cdots,M,$ (15)
@@
$\left\|\mathbf{U}^{H}\mathbf{U}\right\|_{F}^{2}=\left\|\mathbf{U}\mathbf{U}^{H}\right\|_{F}^{2}.$ (17)
@@
$\left\|\mathbf{U}^{H}\mathbf{U}\right\|_{F}^{2}=$ $\sum_{k,k^{\prime}=0}^{2N-2}\left|\sum_{m=1}^{M}\sum_{r=0}^{Z_{y}-1}\sum_{i=0}^{Z_{x}-1}u(\bm{x}^{m,r})_{i,k}\left(u(\bm{x}^{m,r})_{i,k^{\prime}}\right)^{\ast}\right|^{2}$ (19a)
@@
where the inequality (19b) is obtained by removing all terms $k\neq k^{\prime}$ in (19a). Equation (19c) is derived based on the condition (14b) of the weight vector $\mathbf{p}$. The proof of (19d) is similar to [15, Appendix B] and hence omitted here. ∎
@@
$+MN^{2}\sum_{r=0}^{Z_{y}-1}p_{r}^{2}\sum_{i=0}^{Z_{x}-1}w_{i}^{2}+M^{2}\sum_{d=E}^{D}d^{2}\sum_{\begin{subarray}{c}s,t=0\\ l_{s,t,N}=N-d\end{subarray}}^{Z_{x}-1}w_{s}w_{t},$ (20)
@@
$||\mathbf{U}\mathbf{U}^{H}||_{F}^{2}=$ $\sum_{m,m^{\prime}=1}^{M}\sum_{r,r^{\prime}=0}^{Z_{y}-1}\sum_{i,i^{\prime}=0}^{Z_{x}-1}|A_{\bm{x}^{m},\bm{x}^{m^{\prime}}}(\tau_{i,i^{\prime}},\nu_{r,r^{\prime}})|^{2}p_{r}p_{r^{\prime}}w_{i}w_{i^{\prime}}$
$\leq$ $\sum_{\begin{subarray}{c}m,m^{\prime}=1\\ m\neq m^{\prime}\end{subarray}}^{M}\sum_{r,r^{\prime}=0}^{Z_{y}-1}\sum_{i,i^{\prime}=0}^{Z_{x}-1}{\theta}_{c}^{2}p_{r}p_{r^{\prime}}w_{i}w_{i^{\prime}}+\sum_{m=1}^{M}\sum_{r,r^{\prime}=0}^{Z_{y}-1}\sum_{i,i^{\prime}=0}^{Z_{x}-1}{\theta}_{a}^{2}p_{r}p_{r^{\prime}}w_{i}w_{i^{\prime}}$
@@
$-M(M-1)\sum_{d=E}^{D}\left({\theta}_{c}^{2}-d^{2}\right)\sum_{\begin{subarray}{c}s,t=0\\ l_{s,t,N}=N-d\end{subarray}}^{Z_{x}-1}w_{s}w_{t}-M\sum_{d=E}^{D}\left({\theta}_{a}^{2}-d^{2}\right)\sum_{\begin{subarray}{c}s,t=0\\ l_{s,t,N}=N-d\end{subarray}}^{Z_{x}-1}w_{s}w_{t}.$ (22)
@@
${\theta}_{c}^{2}(M-1)\left(1-\mathbf{w}^{T}\sum_{d=E}^{D}\bm{J}^{d}_{Z_{x}}\mathbf{w}\right)+{\theta}_{a}^{2}\left(1-\mathbf{w}^{T}\left(\bm{I}_{Z_{x}}+\sum_{d=E}^{D}\bm{J}^{d}_{Z_{x}}\right)\mathbf{w}\right)$
@@
${{\theta}^{2}_{\text{max}}}\geq\left\{\begin{matrix}{N-1},&M=1;\\ {N},&M>1.\end{matrix}\right.$ (26a)
@@
For $Z_{x}=Z_{y}=N$, according to (II-B) and Lemma 1, we have
@@
The bounds for global AF in (26) is also a special case of the proposed bound (25b) for delay-Doppler LAZ. When $Z_{y}=N$ and $M\geq 2$, the proposed bound (25b) is equal to (26a) by choosing weight vector $\mathbf{w}=[1,0,\cdots,0]^{T}$. When $Z_{y}=N$ and $M=1$, the proposed bound (25b) is equal to (26a) with weight vector $\mathbf{w}=[0.5,0.5,0,\cdots,0]^{T}$.

### v3

# Generalized Arlery-Tan-Rabaste-Levenshtein Lower Bounds on Ambiguity Function and Their Asymptotic Achievability

@@ Abstract
This paper presents generalized Arlery-Tan-Rabaste-Levenshtein lower bounds on the maximum aperiodic ambiguity function (AF) magnitude of unimodular sequences under certain delay-Doppler low ambiguity zones (LAZ). Our core idea is to explore the upper and lower bounds on the Frobenius norm of the weighted auto- and cross-AF matrices by introducing two weight vectors associated with the delay and Doppler shifts, respectively. As a second major contribution, we demonstrate that our derived lower bounds are asymptotically achievable with selected Chu sequence sets by analyzing their maximum auto- and cross- AF magnitudes within certain LAZ.
@@ I Introduction
The maximum magnitude of the AF sidelobes critically influences the target detection and delay-Doppler estimation capabilities of sequences, as the AF peaks of weaker targets may be buried by the sidelobes of strong targets. To achieve reliable detection and estimation, the ideal sequence sets are expected to have zero periodic/aperiodic auto-ambiguity function (AAF) and cross-ambiguity function (CAF) values, except for the AAF peak at the origin of the delay-Doppler plane. However, such sequence sets are impossible due to the limited volume of AFs [4]. In practical applications, a feasible approach is to optimize the local AF sidelobes of sequences over a low ambiguity zone (LAZ) [5] in the delay-Doppler plane, since the maximum delay and Doppler values are generally much smaller than the signal duration and bandwidth, respectively [6, 7, 8, 9, 10, 11].
@@
Correlation functions can be considered as a special case of AFs in the zero-Doppler cut. In the literature, two important correlation lower bounds are the Welch bound [12] and the Levenshtein bound [13, 14, 15]. These bounds are useful measures of sequence sets operating under static (or quasi-static) wireless environments only. In contrast, the upcoming sixth generation (6G) mobile systems need to 1) deal with highly dynamic environments with high mobility of 1000 km/h (or higher) [16, 17] and 2) support efficient design of integrated sensing and communication (ISAC) systems [18, 19]. Driven by these two important applications, it is desirable to understand the AF lower bounds in order to measure the optimality of sequence sets in asynchronous mobile channels. Although periodic AF lower bounds are investigated in [20] and [5], a comprehensive study of the aperiodic AF lower bounds is still lacking. There is an aperiodic AF lower bound in [5] but it is derived by extending the methodologies from the periodic case. Furthermore, the trade-offs between the lower bounds of the aperiodic AAF and CAF are largely untouched, to the best of our knowledge.
@@
From the construction perspective, while sequences with low aperiodic correlation have been investigated with respect to the Welch bound [12] and the Pursley-Sarwate criterion [21] in [22, 23, 24, 25, 26], little is known on systematically constructed sequences having low aperiodic AF sidelobes. The asymptotic behavior of Chu sequence pairs [27] was studied in [22] and [23], showing that the maximum auto-correlation and cross-correlation magnitudes of the primary Chu sequence and its conjugate can asymptotically meet the Welch bound [12]. This observation demonstrates that the aperiodic Welch bound is asymptotically achievable, highlighting the effectiveness of Chu sequences. However, similar evidence supporting the achievability of the aperiodic AF lower bounds remains absent.
In this paper, compared to the bound in [5, Theorem 4], we introduce tighter aperiodic AF lower bounds for unimodular sequences with certain LAZ. Such bounds are called “generalized Arlery-Tan-Rabaste-Levenshtein bounds” as we are motivated by some techniques in [13] and [15] focusing on aperiodic correlation lower bounds of polyphase sequence sets. Our core idea is to explore the upper and lower bounds on the Frobenius norm of the weighted auto- and cross-AF matrices by introducing the delay- and Doppler-related weight vectors as well as several useful structural properties of aperiodic AF.
@@
It is shown that, at zero Doppler, the proposed AF lower bounds include not only the Levenshtein bound [13] and the Arlery-Tan-Rabaste bound [15], but also several other aperiodic correlation bounds, such as the Welch bound [12], the Sarwate bound [28], and the Peng-Fan bound [29]. Interestingly, our derivation reveals that the optimal Doppler weight vector should be the one that evenly weights the Doppler bins within the LAZ, aligning with the cyclic nature of the aperiodic AF in the Doppler dimension. In contrast, due to the non-cyclic characteristics of the aperiodic AF in the delay dimension, it is found that tighter aperiodic AF bounds can be obtained by certain selected delay weight vectors.
@@
Additionally, by analyzing the maximum aperiodic AAF and CAF magnitudes of Chu sequences, we show that certain pairs/sets of Chu sequences are asymptotically order-optimal with respect to our proposed AF bounds over certain LAZ. This analysis substantiates the achievability of these bounds. We then demonstrate the tightness of the proposed bounds through asymptotic analysis and numerical examples.
@@ II-A Definitions
$A_{\bm{x}^{m},\bm{x}^{m^{\prime}}}(\tau,\nu)$
$=$ $\left\{\begin{array}[]{ll}\displaystyle\sum_{t=0}^{N-1-\tau}{x}^{m}_{t}{\left({x}^{m^{\prime}}_{t+\tau}\right)}^{\ast}e^{\frac{j2\pi\nu t}{N}},&0\leq\tau\leq N-1;\\ \displaystyle\sum_{t=0}^{N-1+\tau}{x}^{m}_{t-\tau}{\left({x}^{m^{\prime}}_{t}\right)}^{\ast}e^{\frac{j2\pi\nu t}{N}},&1-N\leq\tau\leq-1,\end{array}\right.$ (3)
@@
We further define the LAZ in the delay-Doppler plane as
@@ II-B Properties of Aperiodic AF
Let $\bm{x}=[x_{0},x_{1},\cdots,x_{N-1}]$, for any $\nu\neq 0$, we have $A_{\bm{x}}(0,\nu)=\sum_{t=0}^{N-1}x_{t}{\left(x_{t}\right)}^{\ast}e^{\frac{j2\pi\nu t}{N}}=\sum_{t=0}^{N-1}e^{\frac{j2\pi\nu t}{N}}=0$. ∎
From the definition of $A_{\bm{x}^{m},\bm{x}^{m^{\prime}}}(\tau,\nu)$, the number of terms in its summation is $N-|\tau|$. By letting $d=N-|\tau|$, $d\in[1,N]$, one can readily show the following lemma:
@@
${\left|A_{\bm{x}^{m},\bm{x}^{m^{\prime}}}(\tau,\nu)\right|}\leq{d},\quad\forall\,\,|\tau|,|\nu|\in[0,N-1].$ (8)
@@ III-A Aperiodic AF Lower Bounds Associated to the LAZ
$\mathbf{U}=\begin{bmatrix}{\mathbf{U}^{1}}\\ {\mathbf{U}^{2}}\\ \vdots\\ {\mathbf{U}^{M}}\end{bmatrix},\;\text{with}\;\mathbf{U}^{m}=\begin{bmatrix}{\bm{u}(\bm{x}^{m,0})}\\ {\bm{u}(\bm{x}^{m,1})}\\ \vdots\\ {\bm{u}(\bm{x}^{m,Z_{y}-1})}\end{bmatrix}\;$ (12)
$\text{for}\;m=1,2,\cdots,M,$
@@
$\left\|\mathbf{U}^{H}\mathbf{U}\right\|_{F}^{2}=\left\|\mathbf{U}\mathbf{U}^{H}\right\|_{F}^{2},$ (14)
where $\left\|\cdot\right\|_{F}$ represents the Frobenius norm. Specifically, for a matrix $\mathbf{Y}\in\mathbb{C}^{m\times n}$, we have $\|\mathbf{Y}\|_{F}=\sqrt{\sum_{i=1}^{m}\sum_{j=1}^{n}|Y_{ij}|^{2}}$, with $Y_{ij}$ denoting the $(i,j)$-th entry of $\mathbf{Y}$.
@@
$\left\|\mathbf{U}^{H}\mathbf{U}\right\|_{F}^{2}$
$=$ $\sum_{k,k^{\prime}=0}^{2N-2}\left|\sum_{m=1}^{M}\sum_{r=0}^{Z_{y}-1}\sum_{i=0}^{Z_{x}-1}u(\bm{x}^{m,r})_{i,k}\left(u(\bm{x}^{m,r})_{i,k^{\prime}}\right)^{\ast}\right|^{2}$ (16a)
@@
where the inequality (16b) is obtained by removing all terms $k\neq k^{\prime}$ in (16). Equation (16c) is derived based on the condition (11b) of the weight vector $\mathbf{p}$. The proof of (16d) is similar to [15, Appendix B] and hence omitted here. ∎
{strip}
@@
$+MN^{2}\sum_{r=0}^{Z_{y}-1}p_{r}^{2}\sum_{i=0}^{Z_{x}-1}w_{i}^{2}$
$+M^{2}\sum_{d=E}^{D}d^{2}\sum_{\begin{subarray}{c}s,t=0\\ l_{s,t,N}=N-d\end{subarray}}^{Z_{x}-1}w_{s}w_{t},$ (17)
@@
$||\mathbf{U}\mathbf{U}^{H}||_{F}^{2}$
$=$ $\sum_{m,m^{\prime}=1}^{M}\sum_{r,r^{\prime}=0}^{Z_{y}-1}\sum_{i,i^{\prime}=0}^{Z_{x}-1}|A_{\bm{x}^{m},\bm{x}^{m^{\prime}}}(i-i^{\prime},r-r^{\prime})|^{2}p_{r}p_{r^{\prime}}w_{i}w_{i^{\prime}}$
$\leq$ $\sum_{\begin{subarray}{c}m,m^{\prime}=1\\ m\neq m^{\prime}\end{subarray}}^{M}\sum_{r,r^{\prime}=0}^{Z_{y}-1}\sum_{i,i^{\prime}=0}^{Z_{x}-1}{\theta}_{c}^{2}p_{r}p_{r^{\prime}}w_{i}w_{i^{\prime}}$
$+\sum_{m=1}^{M}\sum_{r,r^{\prime}=0}^{Z_{y}-1}\sum_{i,i^{\prime}=0}^{Z_{x}-1}{\theta}_{a}^{2}p_{r}p_{r^{\prime}}w_{i}w_{i^{\prime}}$
@@
$-M(M-1)\sum_{d=E}^{D}\left({\theta}_{c}^{2}-d^{2}\right)\!\!\!\sum_{\begin{subarray}{c}s,t=0\\ l_{s,t,N}=N-d\end{subarray}}^{Z_{x}-1}\!\!\!w_{s}w_{t}$
$-M\sum_{d=E}^{D}\left({\theta}_{a}^{2}-d^{2}\right)\!\!\!\sum_{\begin{subarray}{c}s,t=0\\ l_{s,t,N}=N-d\end{subarray}}^{Z_{x}-1}\!\!\!w_{s}w_{t}.$ (19)
@@
${\theta}_{c}^{2}(M-1)\left(1-\mathbf{w}^{T}\sum_{d=E}^{D}\bm{J}^{d}_{Z_{x}}\mathbf{w}\right)$
$+$ ${\theta}_{a}^{2}\left(1-\mathbf{w}^{T}\left(\bm{I}_{Z_{x}}+\sum_{d=E}^{D}\bm{J}^{d}_{Z_{x}}\right)\mathbf{w}\right)$
@@
${{\theta}^{2}_{\text{max}}}\geq\left\{\begin{matrix}{N-1},&M=1;\\ {N},&M>1.\end{matrix}\right.$ (23)
@@
$\sum_{\nu=0}^{N-1}{\left|A_{\bm{x}^{m},\bm{x}^{m^{\prime}}}(\tau,\nu)\right|}^{2}$
$=$ $N\sum_{t=0}^{N-1}{\left|x_{t}^{m}\right|}^{2}{\left|x_{t+\tau}^{m^{\prime}}\right|}^{2}=N\left({N-|\tau|}\right).$ (24)
For $Z_{x}=Z_{y}=N$, according to (III-A) and Lemma 1, we have
@@
The lower bound for global AF in (23) is also a special case of the proposed bound (22b) for delay-Doppler LAZ. When $Z_{y}=N$ and $M\geq 2$, the proposed bound (22b) is equal to (23) by choosing weight vector $\mathbf{w}=[1,0,\cdots,0]^{T}$. When $Z_{y}=N$ and $M=1$, the proposed bound (22b) is equal to (23) with weight vector $\mathbf{w}=[0.5,0.5,0,\cdots,0]^{T}$.